OpenStudy (anonymous):

I have a question that says to "use the half-angle identities to evaluate cos(pi/8) exactly" and i can't solve it past +-√[(1+cos(pi/8))/2] ... anyone?

OpenStudy (anonymous):

i went cos(pi/8) to cos(pi/4/2) btw

OpenStudy (anonymous):

I got you :P \[\cos^2(\frac{\pi}{8})=\frac{1}{2}(1+\cos(\frac{\pi}{4}))\] Cosine is positive since pi/4 is going to give a positive x. So: \[\cos(\frac{\pi}{8})=\frac{\sqrt{2+\sqrt{2}}}{2}\]

OpenStudy (anonymous):

At the end, your answer will look different but I did a little manipulation to make it look nice.

OpenStudy (anonymous):

I'm sorry.... :( I don't follow what you did......

OpenStudy (anonymous):

Use the formula: \[\cos^2(a)=\frac{1}{2}(1+\cos(2a))\] a=pi/8 Can you follow from there?

OpenStudy (anonymous):

No - I don't have that formula. I have cos(x/2)=√[(1+cosx)/2]

OpenStudy (anonymous):

oh hell its the same thing

OpenStudy (anonymous):

Its the same formula. Square your formula cos^2(x/2)=(1+cos(x))/2 Either way. Using your (same) formula you have: \[\cos(\frac{\frac{\pi}{4}}{2})=\sqrt{\frac{1+\cos(\frac{\pi}{4})}{2}}\]

OpenStudy (anonymous):

Then you have: cos(pi/4)=sqrt(2)/2 \[\sqrt{\frac{1+\frac{\sqrt{2}}{2}}{2}}=\sqrt{\frac{\frac{2+\sqrt{2}}{2}}{2}}=\frac{\sqrt{2+\sqrt{2}}}{\sqrt{2}*\sqrt{2}}=\frac{\sqrt{2+\sqrt{2}}}{2}\]

OpenStudy (anonymous):

I can't give two medals can I.

OpenStudy (anonymous):

Haha, no :P But thank you.

OpenStudy (anonymous):

dude you're blowin' my mind right now

OpenStudy (anonymous):

Sorry D: Whats wrong?

OpenStudy (anonymous):

no it's good - i can't even see what you saw until you did it. but the 1 = 2/2, so 1 + r2/2 = 2+r2 all over 2 but

OpenStudy (anonymous):

divide two divide two equals divide 4, and r2 *r2 doesn't equal four in the denom....or?

OpenStudy (anonymous):

jesus, wait....lol

OpenStudy (anonymous):

smh

OpenStudy (anonymous):

Haha, I can work it out in slightly more gruesome detail, but not much more :/

OpenStudy (anonymous):

yeah, sorry i got it.

OpenStudy (anonymous):

Sure?

OpenStudy (anonymous):

100% - i don't know why i didn't apply the radical to the denominator. crap :( much thanks bud

OpenStudy (anonymous):

No problem :P

OpenStudy (anonymous):

chess was easily to master than trig

OpenStudy (anonymous):

easier? yeah

OpenStudy (anonymous):

Haha, I just made that mistake :P Any other questions just post them man :P

OpenStudy (anonymous):

thanks bud. again much thanks.